teddy has simplify 10 cube root 48 here is the answers 10 square root 48=10x4x12=20root12 do you agree explain why



1
Expert's answer
2021-03-05T07:39:36-0500

10 cube root 48 = 10 * (48)(1/3)

= 10 * (2 * 2 * 2 * 6)(1/3)

= 10 * 2 * (6)(1/3)

= 20 * 1.817120 = 36.34



And


20root12 = 20 * (12)(1/2) = 20 * 2 * (3)(1/2) = 40 * 1.73 = 69.28

10 square root 48 = 10 * (48)(1/2) = 10 * (2*2*12)(1/2) = 20*(12)(1/2) = 20 * 3.46 = 69.28


Therefore, 20 root 12 == 10 square root 48